OpenStudy (anonymous):

sqrt-64 / (7-6i) - (2-2i) choices: -64+40i/-3 64+40i/-3 -32+40i/41 32+40i/41

OpenStudy (ash2326):

We have \[\frac{\sqrt {-64}}{(7-6i-(2-2i))}\] Now we know that \(\sqrt {-1}=i\) so \[\frac{\sqrt {64}\times \sqrt {-1}}{(7-6i-(2-2i))}\] \[\frac{8i}{(7-6i-(2-2i))}\] Let's simplify the denominator \[\frac{8i}{(7-6i-2+2i)}\] \[\frac{8i}{(5-4i)}\] Let's multiply the numerator and denominator by 5+4i \[\frac{8i}{(5-4i)}\times \frac{5+4i}{5+4i}\] now we get \[\frac{40i+32i^2}{25+16}\] we get \[\frac{-32+40i}{41}\]